英語 (English)

基盤英語 (Basic English) (総科 (全 1 年))

山田仁子 · 准教授/ 大学院ソシオ・アーツ・アンド・サイエンス研究部 1 単位前期火 3 ・ 4

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

【授業の目的】英語を聞いて理解する力を養う．このためには，英語音声に慣れ

ること，語彙力や文法的知識をを高めることが必要となる．【授業の概要】世界遺産を題材とする DVD 教材を視聴し，まずは大まかな内容

を捉え，次に単語や表現など細部まで理解するようにする．【キーワード】 listening comprehension， vocabulary building 【先行科目】 [先行科目] 【関連科目】 [関連科目] 【到達目標】

1. 英語を聞いて内容を理解できるようになる．

2. 語彙力を高める. 【授業の計画】

1. 授業について説明

2. Chapter 15 The Memory of the Holocaust: Auschwitz (Poland) (日) Chapter 15 The Memory of the Holocaust: Auschwitz (Poland)

3. Chapter 14 Reawakened City: Archaeological Areas of Pompeii (Italy)

4. Chapter 13 A Landscape Shaped by Glaciers: Te Wahipounamu (New Zealand)

5. Chapter 12 The Water Palace: The Alhambra, Granada (Spain)

6. Chapter 11 A Nation of Civility: Gusuku Sites and Related Properties of the Kingdom of Ryukyu (Japan)

7. Chapter 10 Underground Towns: Goreme National Park and the Rock Sites of Cappadocia (Turkey)

8. 中間試験

9. Chapter 9 A Tragic Masquerade: The Royal Palace of Drottningholm (Sweden)

10. Chapter 8 The Maze City: The Medina of Fez (Morocco)

11. Chapter 7 The Challenge to Find Longitude: Greewich (United Kingdom)

12. Chapter 6 Hidden City: Historic Site of Machu Picchu (Peru)

13. Chapter 5 Beautiful Paris, Forever: Paris and the Banks of the Seine (France)

14. Chapter 4 Where Ancient Spirits Live: Uluru-Kata Tjuta National Park (Australia)

15. 期末試験 16. 総括授業【教科書】 DVD でめぐる世界遺産成美堂【参考書等】 [参考資料] 【成績評価の方法】 2 回のテスト，出席状況，予習の程度，授業中の発表などを総合

的に評価する．【再試験の有無】なし．【受講者へのメッセージ】授業開始前までに自習用の DVD を視聴し，教科書の問題を解いてくること. 予習も成績に反映されます. 授業中には積極的に参加する事も期待します.

【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220813 【連絡先 (オフィスアワー・研究室・ E メールアドレス)】

⇒ 山田 (hitokoy@yahoo.co.jp) ���� (オフィスアワー: 木曜日 10:20˜ 11:20)

English

Basic English (総科 (全 1 年))

Hitoko Yamada · Associate Professor / Institute of Socio-Arts and Sciences 1 unit 前期火 3 ・ 4

(平成 19 年度以前の授業科目: 『基盤英語』 ) (平成 16 年度以前 (医保は 17 年度以前) の授業科目: 『英語 (1)』 )

Target⟩ 英語を聞いて理解する力を養う．このためには，英語音声に慣れること，

語彙力や文法的知識をを高めることが必要となる． Outline⟩ 世界遺産を題材とする DVD 教材を視聴し，まずは大まかな内容を捉

え，次に単語や表現など細部まで理解するようにする． �������⟩ listening comprehension, vocabulary building Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Goal⟩

1. 英語を聞いて内容を理解できるようになる．

2. 語彙力を高める. Schedule⟩

1. Introduction

2. Chapter 15 The Memory of the Holocaust: Auschwitz (Poland) ( ) Chapter 15 The Memory of the Holocaust: Auschwitz (Poland)

3. Chapter 14 Reawakened City: Archaeological Areas of Pompeii (Italy)

4. Chapter 13 A Landscape Shaped by Glaciers: Te Wahipounamu (New Zealand)

5. Chapter 12 The Water Palace: The Alhambra, Granada (Spain)

6. Chapter 11 A Nation of Civility: Gusuku Sites and Related Properties of the Kingdom of Ryukyu (Japan)

7. Chapter 10 Underground Towns: Goreme National Park and the Rock Sites of Cappadocia (Turkey)

8. Mid-Term Exam.

9. Chapter 9 A Tragic Masquerade: The Royal Palace of Drottningholm (Sweden)

10. Chapter 8 The Maze City: The Medina of Fez (Morocco)

11. Chapter 7 The Challenge to Find Longitude: Greewich (United Kingdom)

12. Chapter 6 Hidden City: Historic Site of Machu Picchu (Peru)

13. Chapter 5 Beautiful Paris, Forever: Paris and the Banks of the Seine (France)

14. Chapter 4 Where Ancient Spirits Live: Uluru-Kata Tjuta National Park (Australia)

15. Final Exam.

16. 総括授業 Textbook⟩ Exploring World Heritage on DVD Seibido Reference⟩ [参考資料] Evaluation Criteria⟩ 2 回のテスト，出席状況，予習の程度，授業中の発表など

を総合的に評価する． Re-evaluation⟩ なし． Message⟩ 授業開始前までに自習用の DVD を視聴し，教科書の問題を解いてくること. 予習も成績に反映されます. 授業中には積極的に参加する事も期待します.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=220813 Contact (Office-Hour, Room, E-mail)⟩

⇒ Yamada (hitokoy@yahoo.co.jp) ���� (Office Hour: 木曜日 10:20˜ 11:20)